performance of 40 teachers in the Mathematics department showed a mean of 95 with a standard deviation of 7.3.


1
Expert's answer
2021-11-30T11:01:45-0500

The performance of 50 teachers from the faculty of engineering showed a mean of 98 with a standard deviation of 9.2, while, a sample of 40 teachers from the college of science showed an average performance of 95 with a standard deviation of 7.3. Is there a difference in performance between the two samples? Use α = 0.05

The following null and alternative hypotheses need to be tested:

"H_0: \mu_1 = \mu_2 \\

H_1: \mu_1 ≠ \mu_2"

This corresponds to a two-tailed test.

Level of significance =0.05

Sincesample sizes are greater than 30 so z-test for two means, with unknown population standard deviations will be used.

Under null hypothesis the test statistic is obtained as:

"Z = \frac{\bar{x_1} - \bar{x_2}}{\sqrt{s^2_1/n_1 + s^2_2/n_2}} \\

= \frac{98-95}{\sqrt{(9.2)^2/50 + (7.3)^2/40}} \\

= 1.725"

Based on the information provided, the significance level is 0.05, and the critical value for a two-tailed test is "Z_c =1.96"

The rejection region for this two-tailed test is

"R = {Z:|Z|>1.96}"

Since it is observed that

"|Z| = 1.725 < Z_c = 1.96"

It is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2" at 0.05 significance level.